If Kate earns $16 interest in an investment of $425, find the variation equation for I
Answer by stanbon(75887)

If Kate earns $16 interest in an investment of $425,
find the variation equation for I
-----
I = Prt
---
r = I/(Pt)
---
rate = 16/425 = 0.03765 = 3.765%
---------------
Cheers,
Stan H.
---------------